OpenStudy (anonymous):

Is (-4, 2) a solution of 3y - 2x = 14?

OpenStudy (anonymous):

The way you check is to substitute the x and y values from (-4, 2) into the left side of the equation and seeing if it equals the right side(14). 3y - 2x = 14 LS = 3(2) - 2(-4) = 6 + 8 = 14 LS = RS Therefore it is a solution.